# Post-glitch Recovery and the Neutron Star Structure: The Vela Pulsar

### v1

@@ Abstract
Pulsar glitches, sudden changes in a neutron star’s rotation, offer a unique window into the extreme physics governing these celestial bodies. The post-glitch recovery phase, characterized by a gradual recovery, reveals insights into the superfluid dynamics within the star. The post-glitch recovery phase probes the coupling mechanisms between different stellar components, sheds light on the properties of neutron star matter at supra-nuclear densities and even allows us to predict the time to the next glitch. In this work, we present a detailed analysis of the Vela pulsar’s rotational behavior using around 100 months of observational data spanning from September 2016 to January 2025, during which four glitches were identified. Here, we demonstrate the post-glitch recovery of these glitches within the framework of the vortex creep model, providing new insights into the Vela pulsar’s internal structure. Notably, we present the first-ever investigation of vortex residuals – the discrepancy between observed values and those predicted by the vortex creep model – through the lens of the vortex bending model, marking a significant step forward in understanding neutron star physics through pulsar glitches.
I Introduction
@@
Pulsars are highly magnetized rotating dense and compact stars observed to have periodic radiation of pulses with the periodicity tied to its remarkably stable rotation. The slow-down in the rotation of a pulsar is influenced by two types of instabilities: glitches and timing noise (D’alessandro, 1996; Haskell and Melatos, 2015; Zhou et al., 2022). Glitches are abrupt changes in the rotational frequency and spin-down rate of a pulsar, while timing noise is a stochastic wander in the spin parameters of pulsars. The occurrence of glitches, as well as their slow relaxations ranging from a few days to several years, reveal the presence of neutron superfluidity inside the inner crust of the neutron star. The most successful mechanism explaining glitches and their recoveries is based on the superfluid vortex pinning and creep model (Anderson and Itoh, 1975; Alpar et al., 1984a, b). This model suggests that neutron superfluidity in the interior of the star leads to a differential rotation, storing a significant amount of angular momentum. When this angular momentum is released, the crust spins up, resulting in a glitch. Pulsar glitches thus provide valuable insights into the superfluid dynamics within the star (Zhou et al., 2022).
@@
The post-glitch recovery is the process in which the pulsar gradually relaxes back towards its original pre-glitch rotational state. According to the vortex creep model, this process is driven by the repinning of superfluid vortices to the crustal nuclei and the re-establishment of the equilibrium lag between the rotational rates of the superfluid interior and the crust. Hence, the vortex dynamics play a crucial role in determining the recovery behavior. The post-glitch relaxation typically involves the contribution of angular momentum transfer from the neutron star interior to the surface, seen as an exponential recovery followed by a linear decay in the spin-down rate. Occasionally, some inward movement of vortices may also play a role (Akbal et al., 2015) where crustquakes initiate glitches. The linear type of relaxation in time originates from the tangling or bending of the vortices in the non-linear creep regions (Gügercinoğlu et al., 2023). The post-glitch recovery can be utilized to probe the internal dynamics and structure of neutron stars (Gügercinoğlu et al., 2022).
@@
PSR J0835–4510, B0833–45 or the Vela pulsar, has shown 26 glitches, most of which are large glitches and detected after a regular, quasi-periodic inter-glitch time. The Vela pulsar is one of the most studied glitching pulsars. The first-ever glitch detection was recorded in the Vela pulsar (Radhakrishnan and Manchester, 1969; Reichley and Downs, 1969). Moreover, the first-ever glitch detection caught in the act was also realized for the Vela pulsar (Palfreyman et al., 2018). The Vela pulsar is titled the most popular or most studied glitching pulsar among other glitching pulsars for various reasons: it is a bright source, young, and displays glitches at regular intervals. The recoveries of glitches also demonstrate strikingly similar behavior. A measure of a pulsar’s slow down over time is described by the braking index (). It is defined by a power-law relation between the pulsar’s spin-down rate and its rotational angular frequency (), given by . The braking index provides insights into the pulsar’s rotational evolution, magnetic fields, and emission mechanisms.
@@
Here, we present the observational behavior of the Vela pulsar from September 2016 to January 2025; during this period, four large glitches were observed in our monitoring campaign (Grover et al., 2024a; Basu et al., 2020). Theoretical inferences from the post-glitch recovery phase are also presented. We have utilized the recovery behavior of four glitches in the Vela pulsar to estimate the following parameters: (a) the fractional moments of inertia (FMI) of various layers participating in a glitch, (b) the re-coupling timescale of the crustal superfluid, (c) theoretical prediction for the time to the next glitch, (d) braking index of the Vela pulsar. The paper’s outline is as follows: we describe the pulsar observations, the data reduction and the analysis processes in section II. A brief description of the vortex creep and bending models is given in Section III. In Section IV, we discuss the results. The conclusions and future work are given in Section V.
II Observations and Analysis
II.1 Sample and Observational Facilities
We use two large radio telescopes to monitor pulsars: the upgraded Giant Metrewave Radio Telescope (uGMRT)(Gupta et al., 2017) and the Ooty Radio Telescope (ORT) (Swarup et al., 1971). Observations of PSR B083345 were part of a regular glitch monitoring program, where we regularly monitor 24 pulsars (i) to detect glitches, (ii) model timing noise, and (iii) model post-glitch recovery. This work presents a comprehensive analysis of the Vela pulsar’s rotational evolution, spanning from September 2016 to January 2025, as part of our ongoing monitoring campaign (Grover et al., 2024a; Basu et al., 2020). Our study primarily utilizes observations from the ORT, supplemented by uGMRT and archival ATNF data to fill temporal gaps and facilitate specialized investigations.
The ORT (Swarup et al., 1971) is a 530m 30m parabolic cylindrical reflector. It is located on a hill with the same slope as the geographic latitude (11°), allowing it to track sources for about 8 hours and covers a declination range of -60° to +60°. Pulsar observations at the ORT are carried out at the central observation frequency of 326.5 MHz using the Pulsar Ooty Radio Telescope New Digital Efficient Receiver, PONDER (Naidu et al., 2015), which produces real-time coherent dedispersed time series data. The typical cadence of the observations at the ORT is 1–14 days.
The GMRT is an interferometer with 30 parabolic antennas, with a diameter of 45 meters each (Swarup et al., 1991; Ananthakrishnan, 1995). The uGMRT (Gupta et al., 2017) covers a declination range of -53°to +90°and supports observations from 150 MHz to 1460 MHz in four bands. The Vela pulsar is observed in uGMRT Band 3 (250–500 MHz) and Band 4 (550–950 MHz). The central square antennas with a subset of nearest arm antennas were used in a phased array synthesizing a large single dish for observations. The outer arm antennas were not included in the phased array due to significant drift in the phase of the voltage outputs, caused by ionospheric variations, in these antennas. A digital filterbank, implemented in the GMRT correlator, was used to record either 1024 or 2048 channel across the bandpass to obtain frequency resolved data with a typical sampling time of 655 s. The cadence of the observations is 15-30 days. We utilize the frequency-resolved uGMRT data primarily to examine the impact of dispersion measure (DM) offsets on the spin-down rate. Additionally, uGMRT data filled a temporal gap around epoch 60300, providing crucial estimates not covered by ORT due to maintenance downtime.
The Commonwealth Scientific and Industrial Research Organisation (CSIRO) Data Access Portal is used to search and access the Parkes pulsar archival observations111https://data.csiro.au/domain/atnf (Hobbs et al., 2011) of the Vela pulsar. These data were mainly used to fill the temporal gaps in glitch number 2, or MJD 58515 glitch, caused by significant observational gaps at ORT. We directly downloaded and utilized the reduced files (or PSRFITS (Hotan et al., 2004) files) and performed subsequent analysis, as described in the next subsection.
II.2 Data Reduction and Analysis
We utilized the DSPSR software (van Straten and Bailes, 2011) and the PINTA pipeline (Susobhanan et al., 2021) to reduce the raw data of the Vela pulsar observed at the ORT and the uGMRT, respectively, into the PSRFITS files with the ephemeris created in our initial observations. The initial analysis, including the generation of a noise-free template and Time of Arrivals (ToAs), was performed using PSRCHIVE (Hotan et al., 2004; van Straten et al., 2012). The ToAs were obtained through the cross-correlation of the template profile with all other observed profiles using the ‘pat’ command, utilizing the Fourier Phase Gradient (PGS) method (Taylor, 1992). The precision timing analysis of the pulsar was done with the help of the pulsar timing software TEMPO2 (Hobbs et al., 2006; Edwards et al., 2006). The DM offsets were estimated using the DMcalc (Krishnakumar et al., 2021) script from the frequency-resolved data from uGMRT.
The rotational evolution of a pulsar during a glitch can be expressed as a Taylor series expansion,
where is the spin frequency, and and are the first and second derivatives of the frequency, respectively, at reference epoch . The spin evolution of the Vela pulsar from September 2016 to January 2025 obtained after precision timing is given in Fig. 1. Each measurement is made with the local timing of several epochs, typically within a 50-200 day data window, contingent on data availability. Exceptions include the first and last few points adjacent to glitches. The resulting measurements are reflected in the precision of error bars, which are smaller than the symbol used in the figure. By choosing multi-epoch windows, we achieve reliable rotational estimates with minimized error bars, providing a robust foundation for studying vortex residuals. This approach effectively averages out temporal effects caused by physical processes, such as ionized interstellar medium scattering-induced broadening, thereby validating the accuracy of measurements aligned with the star’s rotation and internal dynamics.
For each recovery, we utilized Bayesian statistics for model comparison and parameter estimation. Bayesian statistics uses Bayes’ theorem to enhance the probability of the hypothesis based on new evidence or data. For model comparison, the standard Jeffreys scale is utilized, which suggests the preferred model based on the strength of evidence (Trotta, 2008). Bayesian inference is also used to predict the epoch of the next glitch using data from previous glitches.
The priors are selected based on two criteria: initial estimates from the weighted least-square fit and reasonable ranges for each parameter using literature knowledge. These reasonable ranges can be fixed for some parameters, like the fractional moment of inertia (typically positive and below 1) and may vary for others like the recoupling timescale, usually depends whether it is a 3 or 2 or 1 exponential component model. For the linear relaxation region, the prior for recoupling timescale is 10-300 days, and the offset time is 300 to 1200 or more for some recoveries with the initial guess as given by the weighted least-square technique. All the priors were considered as uniform priors. Since all the recoveries are unique hence this flexible approach to setting priors is necessary. The Gaussian likelihood is considered for the analysis. The MCMC Hammer: emcee (Foreman-Mackey et al., 2013) ensemble sampler is used for both parameter estimation and model comparison.
The post-glitch recoveries are studied within the framework of the vortex creep model and the vortex bending model (Gügercinoğlu et al., 2023). We first model the frequency derivative measurements from the data using the vortex creep model. Then, we estimate the residuals obtained after subtracting the predicted value by the vortex creep model from the measured values. These residuals were investigated within the framework of the vortex bending model. The generalized Lomb-Scargle periodogram technique (Lomb, 1976; Scargle, 1982; VanderPlas, 2018) was employed to detect periodicities in the residuals. Peaks with false alarm probabilities (FAP) below 5% or Z-score above 3 were considered significant for vortex bending studies. FAP estimation was performed using the Astropy library (Astropy Collaboration et al., 2013), which implements the Baluev (2008) approach. Alternative methods, including approximations based on effective independent frequencies and bootstrap resampling, yielded similar results. The Z-score significance is obtained from the tail of the one-sided Gaussian distribution corresponding to the observed FAP (Cowan et al., 2011; Ganguly and Desai, 2017). The formalism of the vortex creep and bending models is described in the next section.
III The Vortex Creep and Bending Models
The Vortex Creep model describes the occurrence of glitches and the post-glitch recovery as a byproduct of the superfluid present inside the star. The important features of vortex creep for this paper are now described.
@@ III The Vortex Creep and Bending Models
In the neutron star crust, vortex lines coexist with lattice nuclei, and it is energetically favorable for the vortices to pin to the crystal lattice nuclei. The distribution of vortices is non-uniform due to crustal inhomogeneities such as dislocations, cracks, and impurities formed after quakes. This results in formation of structures such that regions of high vortex density are surrounded by areas with vortex depletion regions collectively referred to as vortex traps (Cheng et al., 1988). Glitches may arise due to collective vortex unpinning events induced in these vortex traps when the angular velocity lag between the superfluid and the crustal normal matter exceeds a threshold value.
After a glitch, the steady-state lag decreases, causing the cessation of creep in certain superfluid regions. The post-glitch recovery process involves the restoration of the creep until the steady-state conditions are re-established by the external braking torque. The vortex creep model describes the formation, recovery and other phenomena related to glitches as follows (Alpar et al., 1984a; Gügercinoğlu and Alpar, 2020; Gügercinoğlu et al., 2022):
A crustquake initiates the formation of vortex traps and triggers a vortex unpinning avalanche. The unpinned vortices transfer their excess angular momentum to the crust, causing it to spin up, which is observed as a glitch.
Superfluid regions in which unpinning occurs decouple from the external decelerating torque, resulting in a step increase in the spin-down rate.
Soon after the large-scale vortex unpinning avalanche, the excess angular momentum of the crustal superfluid is first transferred to the normal matter nuclei and electrons in the crust. This transfer causes a peak in the crustal rotation rate following a glitch.
This change in the crustal rotational state is then communicated to the core superfluid via electrons scattering off magnetized vortex lines (Alpar et al., 1984b). Consequently, the initial peak in the observed crustal rotation rate relaxes on the crust-core coupling time scale, eventually reaching the post-glitch equilibrium rotation rate.
A peak in the rotation rate of the normal component will occur if the coupling between the crustal superfluid and the normal matter is stronger than the coupling between the crustal superfluid and the core superfluid (Pizzochero et al., 2020).
The slow post-glitch evolution is the response of the crustal superfluid to the glitch-induced changes. On longer time scales, the core superfluid reaches equilibrium with the normal matter in the crust, and they behave as a single component. Consequently, the only component decoupled from the braking torque is the crustal superfluid.
The crust breaking can produce transient magnetospheric signatures that coincide with the glitch, such as changes in pulse shape, polarization, and emission. These changes are linked to the plastic motion of the magnetic field lines anchored to the fractured crustal platelets.
Damped sinusoidal-like oscillations in the spin-down rate can appear due to the bending of the vortex creep lines in the linear creep relaxation region, which is mainly generated by the shocks from the respective glitch and can also have an impact from the previous glitches (Gügercinoğlu et al., 2023).
@@
The post-glitch spin-down rate behavior mirrors the responses of the vortex creep regions in both linear and non-linear regimes to a glitch, with additional involvement of induced inward vortex motion in delayed spin-up events. In certain regions of the superfluid in the inner crust and the outer core, vortex creep exhibits a linear response to glitch-induced changes, leading to exponential relaxation (Alpar et al., 1989; Flanagan, 1990; Gügercinoğlu, 2017) as:
@@
where , is the spin frequency, is the spin frequency just before the glitch, is the FMI of the of ith layer that participated in a glitch, and is the exponential recovery timescale.
The exponential decay time scale in the linear creep region of the crustal superfluid relates with the temperature , pinning energy between the vortex line and nucleus , the critical lag for unpinning , the superfluid angular velocity , the Boltzmann constant , and the distance of vortex lines from the rotation axis as (Alpar et al., 1989):
where the microscopic vortex velocity around nuclei, m s-1 throughout the crust (Gügercinoğlu and Alpar, 2016).
The post-glitch response in the non-linear regime for a single crustal superfluid region is (Alpar et al., 1984a):
where is the FMI of the single crustal superfluid region in the non-linear regime, is the pre-glitch spin-down rate, is the recoupling timescale for the single crustal superfluid region in the non-linear regime and the offset time is .
@@
The overall linear relaxation can be obtained by integrating the contributions of several neighboring single non-linear regime regions, i.e., equation 4 with an assumption of linearly decreasing superfluid angular velocity during a glitch (Alpar et al., 1996) and is given as:
where is the FMI of the non-linear crustal superfluid region, is the offset time, and is the non-linear creep relaxation time.
The recoupling timescale in the non-linear creep region of the crustal superfluid relates with the temperature , pinning energy between the vortex line and nucleus , the critical lag for unpinning , the superfluid rotational velocity rate , and the Boltzmann constant as (Alpar et al., 1984a):
@@
The non-linear creep response gives an estimate for the time to the next glitch. The glitch-induced changes in the spin-down rate fully relax after a waiting time (Gügercinoğlu et al., 2022),
@@
where is the observed glitch amplitude and represent the decrement in the superfluid rotational velocity. The post-glitch recovery completes once the relaxation of the non-linear creep regions is finished. This gives us a theoretical prediction for the time to the next glitch in terms of the vortex creep model, which lies between the range to .
An external effect like a crustquake can add an inward vortex motion. This increases the superfluid rotation rate as the vortices pinned to the broken plate are induced to move towards the rotation axis along with the plate. When vortices move inward, the superfluid rotates faster. The lag increases from its steady-state value, making the creep more efficient than in steady state and resulting in an enhanced vortex creep current in the radially outward direction (Gügercinoğlu and Alpar, 2019). The post-glitch response for inward moving vortices causing delayed spin-up observed from some glitches is given as (Akbal et al., 2015):
where is the FMI of the non-linear creep superfluid region in which inward vortex motion took place, and and are corresponding recoupling timescale and offset time associated with inward vortex motion.
@@
A typical post-glitch recovery involves an exponential and a linear relaxation described by equations 2 and 5, respectively. A post-glitch recovery where delayed spin-ups have been observed requires an additional non-linear regime response term provided by equation 9. One or more of these is relevant to the post-glitch behaviour of the Vela pulsar investigated in this paper. This is tested by using different alternative models as described below:
@@
Model 3: Exponential and linear recoveries, the sum of equations 2 and 5.
@@
The evidence for these different alternatives was evaluated using Bayesian analysis of our post-glitch recovery measurements of the Vela pulsar to select the suitable model and estimate its parameters.
Apart from long-term post-glitch recovery, we see evidence for sinusoidal-like spin-down rate oscillations over and above the recovery predicted by the vortex creep model as described in Section IV. Such oscillations can be explained by vortex bending (Gügercinoğlu et al., 2023) and appear to be a common feature following glitches, as also observed from PSRs J1522–5735 (Zhou et al., 2024), J0742–2822 (Zubieta et al., 2025a), J1509–5850 and J1718–3825 (Liu et al., 2025). To understand this better, we used three alternative hypotheses as described below

### v2

@@ Abstract
We present a detailed analysis of the Vela pulsar’s rotational behaviour using approximately 100 months of observational data spanning from September 2016 to January 2025, during which four glitches were identified. Here, we demonstrate the post-glitch recovery of these glitches within the framework of the vortex creep model. We further present the investigation of vortex residuals (the discrepancy between observed values and those predicted by the vortex creep model) by interpreting them in the context of the vortex bending model. In addition, we report a positive correlation between the glitch magnitude and the time to the next glitch, applicable only for the large glitch events observed in the Vela pulsar. Furthermore, we estimate the braking index of the Vela pulsar to be 2.94 0.55.
keywords:
radio astronomy; pulsars: Vela; astronomy data analysis; pulsar timing method;doi:
xx.yyyy/pasa.zzzz.tt1 Introduction
@@
Pulsars are highly magnetised, rotating, dense, and compact stars observed to have periodic radiation of pulses with the periodicity tied to their remarkably stable rotation. The slowdown in the rotation of a pulsar is influenced by two types of instabilities: glitches and timing noise (D'alessandro96; Haskell_2015; zhou_erbil_glitch_review). Glitches are abrupt changes in the rotational frequency and spin-down rate of a pulsar, while timing noise is a stochastic wander in the spin parameters of pulsars. The occurrence of glitches, as well as their slow relaxations, which range from a few days to several years, reveals the presence of neutron superfluidity within the inner crust of the neutron star. One of the most widely accepted mechanisms for explaining glitches and their recoveries is the superfluid vortex pinning and creep model (AndersonItoh1975; Alpar1984I; Alpar1984II). This model suggests that neutron superfluidity in the interior of the star leads to a differential rotation, storing a significant amount of angular momentum. When this angular momentum is released, the crust spins up, resulting in a glitch. Pulsar glitches thus provide valuable insights into the superfluid dynamics within the star (zhou_erbil_glitch_review; Liu2024).
@@
The post-glitch recovery is the process in which the pulsar gradually relaxes back towards its original pre-glitch rotational state. According to the vortex creep model, this process is driven by the repinning of superfluid vortices to the crustal nuclei and the re-establishment of the equilibrium lag between the rotational rates of the superfluid interior and the crust. The specific timescale and form of the relaxation, whether exponential or linear, are determined by the distinct vortex dynamics at play. While the standard recovery reflects the transfer of angular momentum from the interior to the surface, deviations from this norm can indicate complex interactions, such as crustal events or non-linear creep regimes (Akbal_2015; Erbil_2023). Ultimately, the post-glitch recovery serves as a probe for the internal dynamics and structure of neutron stars (Erbil_2022).
@@
PSR J0835–4510, B0833–45, the Vela pulsar, has 26 reported glitch events, most of which are large glitches and detected after a regular, quasi-periodic inter-glitch time111http://www.jb.man.ac.uk/pulsar/glitches.html (Espinoza2011; Basu2022). The Vela pulsar is one of the most studied glitching pulsars. The first-ever glitch was recorded in the Vela pulsar (Radhakrishnan_Manchester_1969; Reichley_Downs_1969). Moreover, the first-ever glitch detection caught in the act was also realised for the Vela pulsar (Palfreyman2018_Vela). The Vela pulsar is considered one of the most popular glitching pulsars for several reasons: it is a bright, young source that displays glitches at regular intervals. The recoveries of glitches also demonstrate strikingly similar behaviour. A measure of a pulsar’s slowdown over time is described by the braking index (). It is defined by a power-law relation between the pulsar’s spin-down rate and its rotational angular frequency (), given by . The braking index provides insights into the pulsar’s rotational evolution, magnetic fields, and emission mechanisms.
@@
Here, we present the observational behaviour of the Vela pulsar from September 2016 to January 2025; during this period, four large glitches were observed in our monitoring campaign (Grover2024; Basu2020). Theoretical inferences from the post-glitch recovery phase are also presented. We have used the recovery behaviour of four glitches in the Vela pulsar to estimate the following parameters: (a) the fractional moments of inertia (FMI) of various layers participating in a glitch, (b) the re-coupling timescale of the crustal superfluid, (c) theoretical prediction for the time to the next glitch, (d) braking index of the Vela pulsar. The paper’s outline is as follows: a brief description of the vortex creep and bending models is given in Section 2. We describe the pulsar observations, the data reduction, and the analysis processes in section 3. In Section 4, we present the results. The conclusions and future scope are given in Section 5.
2 The Vortex Creep and Bending Models
@@
The vortex creep model provides a microscopic framework for understanding the macroscopic behaviour of glitches and their subsequent recovery. In the neutron star crust, vortex lines coexist with and energetically pin to the crystal lattice. Crustal inhomogeneities such as dislocations, cracks, and impurities produced by quakes create a non-uniform vortex distribution known as vortex traps (Cheng1988). Glitches may arise due to collective vortex unpinning events induced in these vortex traps when the angular velocity lag between the superfluid and the crustal normal matter exceeds a threshold value (AndersonItoh1975; Alpar1984I). During the unpinning events, angular momentum is rapidly transferred from the superfluid to the crust. The superfluid regions involved in the unpinning decouple from the external braking torque, resulting in a step increase in the spin-down rate.
After a glitch, the steady-state lag decreases, temporarily suppressing creep in some superfluid regions. The recovery process involves restoring the creep until steady-state conditions are re-established by the external braking torque. Shortly after the glitch, excess angular momentum is transferred from the crustal superfluid to the normal matter in the crust, producing a transient peak in the crustal rotation rate. This perturbation is subsequently communicated to the core superfluid through electron scattering off magnetised vortex lines (Alpar1984II), causing the initial peak to relax on the crust–core coupling timescale toward a new equilibrium state. A pronounced peak in the rotation rate of the normal component is expected when the coupling between the crustal superfluid and the normal matter is stronger than that with the core superfluid (Pizzochero2020). On longer timescales, the core superfluid recouples, leaving only the crustal superfluid effectively decoupled from the external braking torque.
In some glitches of the Crab pulsar, delayed spin-ups are observed (Wong2001; Shaw2018; Basu2020; Ge_2020). Within the vortex creep framework, these features are attributed to vortex unpinning accompanied by inward vortex motion driven by fractured crustal platelets (Akbal_2015; Erbil_2019). Additionally, damped, sinusoidal-like oscillations in the spin-down rate can arise from vortex-line bending in the non-linear creep relaxation regime, excited by the glitch and potentially influenced by preceding events (Erbil_2023). It is important to note that the vortex creep model is not the unique explanation; several equally viable models successfully explain glitch-triggering mechanisms, delayed spin-ups, and post-glitch relaxation. For a comprehensive discussion, see Haskell_2015; Antonopoulou2022; Antonelli2022; zhou_erbil_glitch_review.
@@
The post-glitch spin-down rate behaviour mirrors the responses of the vortex creep regions in both linear and non-linear regimes to a glitch, with additional involvement of induced inward vortex motion in delayed spin-up events. In certain regions of the superfluid in the inner crust and the outer core, vortex creep exhibits a linear response to glitch-induced changes, leading to exponential relaxation (Alpar1989; Flanagan1990; Erbil_2017aug) as:
@@
where , is the spin frequency, is the spin frequency just before the glitch, is the FMI of the of ith layer that participated in a glitch, and is the corresponding exponential recovery timescale.
@@
The overall linear relaxation, obtained by integrating the contributions of several neighbouring single non-linear regime regions (Alpar1984I; Erbil_2022), with an assumption of linearly decreasing superfluid angular velocity during a glitch (Alpar1996) and is given as:
@@
where is the FMI of the non-linear crustal superfluid region, is the offset time, and is the non-linear creep relaxation time. The non-linear creep response gives an estimate for the time to the next glitch. The glitch-induced changes in the spin-down rate fully relax after a waiting time (Erbil_2022),
@@
where is the observed glitch magnitude and represents the decrement in the superfluid rotational frequency. The post-glitch recovery completes once the relaxation of the non-linear creep regions is finished. This gives us a theoretical prediction for the time to the next glitch in terms of the vortex creep model, which lies between the range to .
@@
A typical post-glitch recovery involves an exponential and a linear relaxation described by equations 1 and 2, respectively. This is tested by using different alternative models as described below:
@@
Model 3: Exponential and linear recoveries, the sum of equations 1 and 2.
@@
The most preferred model and corresponding parameters among the various alternatives were evaluated using Bayesian analysis of our post-glitch recovery measurements of the Vela pulsar. Further details of this analysis are presented in the next section.
Apart from long-term post-glitch recovery, we see evidence for sinusoidal-like spin-down rate oscillations over and above the recovery predicted by the vortex creep model as described in Section 4. Such oscillations can be explained by vortex bending (Erbil_2023) and appear to be a common feature following glitches, as also observed from PSRs J1522–5735 (zhou24), J0742–2822 (zubieta25), J1509–5850, and J1718–3825 (liu25). To understand this better, we used three alternative hypotheses as described below